I need help translating the problem situation to a system of equations.
Sarah Comar's Candy Store sold a total of 53 pounds of jelly beans, selling two kinds of jelly beans. The first kind was priced at $4.45 pound, and the second kind was priced at $1.12 per pound. In all, $125.96 was taken in for the two types of jelly beans. How many pounds of each kind were sold? (Let x represent the number of pounds of the first kind and y represent the number of pounds of the second.

Answer by Theo(13342)

Your system of equations would be as follows:

x = the number of pounds of the first kind of jelly beans.
y = the number of pounds of the second kind of jelly beans.

Total of 53 pounds of jelly beans was sold.

Equation 1 would be:

x + y = 53

The first kind of jelly beans was priced at $4.45 per pound.
The second kind of jelly beans was priced at $1.12 per pound.
The total amount of money made was equal to $125.96

Equation 2 would be:

4.45 * x + 1.12 * y = 125.96

The two equations you have to solve simultaneously are:

x + y = 53 (equation 1)
4.45 * x + 1.12 * y = 125.96 (equation 2)

You can solve by substitution or by elimination.

Either way will get you the same answer.